There are five consecutive odd integers. The sum of the frist two numbers is thirty-five less than the sum of the last three numbers. What are the integers?
Answer by sunny1992(35)

note: if first odd number is 1 than the 2nd will be 3 (that is 1+2) and third odd number will be (1+4) 5 and so on
with ref. to above example suppose five consecutive odd integers are
1st x
2nd x+2
3rd x+4
4th x+6
5th x+8
x+x+2=(x+4+x+6+x+8)-35
2x+2=3x+18-35
2-18+35=3x-2x
x=19